# Majority voting is not good for heaven or hell, with mirrored performance

### v4

@@ 3.1 Relationship between gains under opposite proposals
The following lemma establishes a connection between the implemented capital gains under proposal generators ${\mathbf{{X}}}={\mathbf{{X}}}^{\mu}$ and $-{\mathbf{{X}}}$ (such generators will be called opposite ones) when a voting body is antisymmetric.
@@
For any gain generator ${X}$ with mean $\mu$ and any antisymmetric voting body, it holds that
@@
Proof of Lemma 1.
For an arbitrary given ${\mathbf{{X}}}\in{\mathcal{X}}^{\mu},$ we define ${\mathbf{{Y\!}}_{\mathbf{{X}}}}={\mathbf{{X}}}\hskip 1.1ptI({\mathbf{{X}}})$.
Since the voting body is antisymmetric, it follows from (1) that
${\mathbf{{Y\!}}_{\mathbf{{X}}}}-{\mathbf{{Y\!}}_{-{\mathbf{{X}}}}}={\mathbf{{X}}}\hskip 1.1ptI({\mathbf{{X}}})-(-{\mathbf{{X}}})I(-{\mathbf{{X}}})={\mathbf{{X}}}(I({\mathbf{{X}}})+I(-{\mathbf{{X}}}))={\mathbf{{X}}}.$ (2)
By the definition of ${\mathbf{D}}_{\mathbf{{X}}},$ for any $\omega\in\Omega,$ we have
${\mathbb{E}}\bigl{(}{\mathbf{D}}_{\mathbf{{X}}}(\omega)\,|\>\omega\bigr{)}={{\mathbf{{X}}}(\omega)}\hskip 1.1ptI({{\mathbf{{X}}}(\omega)})+0\cdot\!(1-I({{\mathbf{{X}}}(\omega)}))={\mathbf{{Y\!}}_{\mathbf{{X}}}}(\omega).$
The law of total expectation yields
${\mathbb{E}}({\mathbf{D}}_{\mathbf{{X}}})={\mathbb{E}}({\mathbb{E}}({\mathbf{D}}_{\mathbf{{X}}}(\omega)\,|\>\omega))={\mathbb{E}}({\mathbf{{Y\!}}_{\mathbf{{X}}}}).$ (3)
By (3) and (2), it holds that
${\mathbb{E}}({\mathbf{D}}_{\mathbf{{X}}})-{\mathbb{E}}({\mathbf{D}}_{-{\mathbf{{X}}}})={\mathbb{E}}({\mathbf{{Y\!}}_{\mathbf{{X}}}})-{\mathbb{E}}({\mathbf{{Y\!}}_{-{\mathbf{{X}}}}})={\mathbb{E}}({\mathbf{{Y\!}}_{\mathbf{{X}}}}-{\mathbf{{Y\!}}_{-{\mathbf{{X}}}}})={\mathbb{E}}{\mathbf{{X}}}=\mu\bm{1}.$
@@
The decision-making process with generator ${\mathbf{{X}}}$, viewed in the reverse order, shares the generator with the forward process with $-{\mathbf{{X}}}$. However, under the antisymmetry of the voting body, these processes are complementary in terms of decisions made. This observation, due to the Reviewer, allows us to state that the key antisymmetry condition is equivalent to the requirement that the forward and corresponding backward transitions are evaluated by the voting body oppositely in the sense of Eq. (1).
One more interesting observation by the Reviewer is that Lemma 1 still holds for another model where proposals are generated in the same way, but agents’ votes on proposals (which are changes to the state of the world) stem from their preferences over the states of the world themselves (cf. Subsection 1.3). This origin of votes (or social decisions) along with appropriate tie-breaking rules determine the reversal symmetry of the agents’ strategies (respectively, of the voting body). When it is satisfied, state-based preferences are somewhat more flexible than proposal-based votes.
@@ 3.2 Complementary strategies
Consider an alternative group strategy: the group decides whether to support a proposal using the within-group majority rule. More specifically, the group members vote as individuals; let $n_{G}^{+}({\mathbf{{x}}})$ be the total number of votes they cast for proposal ${\mathbf{{x}}}$ in the internal group voting; finally, each group member $i$ casts $n_{i}^{+}({\mathbf{{x}}})$ votes defined by
$n_{i}^{+}({\mathbf{{x}}})=\begin{cases}1,&n_{G}^{+}({\mathbf{{x}}})>\tfrac{g}{2}\\ \tfrac{1}{2},&n_{G}^{+}({\mathbf{{x}}})=\tfrac{g}{2}\\ 0,&n_{G}^{+}({\mathbf{{x}}})<\tfrac{g}{2}\end{cases}.$ (4)
@@
We say that the strategy of agent $i$ is complementary (or satisfies reversal symmetry) if and only if
$n_{i}^{+}({\mathbf{{x}}})+n_{i}^{+}(-{\mathbf{{x}}})=1\mbox{~{}~{}for any proposal~{}~{}}{\mathbf{{x}}}\in{\mathbb{R}}^{n}.$ (5)
@@
The strategies of individuals and members of utilitarian and majoritarian groups are complementary.
Proof of Lemma 2.
Obviously, for a member of a utilitarian group, it holds that $n_{i}^{+}({\mathbf{{x}}})=\tfrac{1}{2}\bigl{(}1+\mathop{\rm sgn}\nolimits\sum_{j=1}^{g}{{x}_{j}}\bigr{)}$, while for a member of a majoritarian group, we have $n_{i}^{+}({\mathbf{{x}}})=\tfrac{1}{2}\bigl{(}1+\mathop{\rm sgn}\nolimits\sum_{j=1}^{g}\mathop{\rm sgn}\nolimits{{x}_{j}}\bigr{)}.$ It is easy to see that in both cases, Eq. (5) is valid. For an individual, this follows as a special case. ∎
@@ 3.3 A class of antisymmetric voting bodies
The symmetrized majority rule has a form similar to (4):
$I({\mathbf{{x}}})\;=\;\begin{cases}1,&n^{+}({\mathbf{{x}}})>\tfrac{n}{2}\\ \tfrac{1}{2},&n^{+}({\mathbf{{x}}})=\tfrac{n}{2}\\ 0,&n^{+}({\mathbf{{x}}})<\tfrac{n}{2}\end{cases}.$
This rule satisfies the following condition:
[…]
Let us say that a stochastic voting rule is complementary iff it satisfies (6).
@@
Any voting body formed by agents with complementary strategies and a complementary voting rule is antisymmetric.
Proof of Lemma 3.
@@
$n^{+}({\mathbf{{x}}})+n^{+}(-{\mathbf{{x}}})=\sum_{i=1}^{n}n_{i}^{+}({\mathbf{{x}}})+\sum_{i=1}^{n}n_{i}^{+}(-{\mathbf{{x}}})=\sum_{i=1}^{n}\bigl{(}n_{i}^{+}({\mathbf{{x}}})+n_{i}^{+}(-{\mathbf{{x}}})\bigr{)}=n.$
Since the voting rule is complementary, this implies that $I({\mathbf{{x}}})+I(-{\mathbf{{x}}})=1$ for all ${\mathbf{{x}}}\in{\mathbb{R}}^{n}$. Thus, the voting body is antisymmetric.∎
@@
Another example of a complementary voting rule is shown in Fig. 2.
Lemma 3 can be applied to societies that may contain several groups and employ the symmetrized majority rule.
@@
The conditions of antisymmetry of a voting body provided by Lemma 3 are not necessary. Indeed, let the agents support all proposals under the voting rule
$I({\mathbf{{x}}})=\begin{cases}\tfrac{1}{2},&n^{+}({\mathbf{{x}}})=n\\ 0,&n^{+}({\mathbf{{x}}})<n\end{cases}.$
Then the voting body is antisymmetric, while neither the agents’ strategy nor the voting rule are complementary. For this voting body, the conclusion of Lemma 1 takes the form ${\mathbb{E}}({\mathbf{D}}_{-{\mathbf{{X}}}})-{\mathbb{E}}({\mathbf{D}}_{\mathbf{{X}}})=-\tfrac{1}{2}\mu\bm{1}-\tfrac{1}{2}\mu\bm{1}=-\mu\bm{1}$.
@@ 3.4 Complementary strategies and rules on opposite proposals
For any gain generator ${X}$ with mean $\mu$ and for any voting body formed by agents with complementary strategies and a complementary voting rule, it holds that
@@
The following corollary is a consequence of Theorem 5, Lemma 2 and the fact that the symmetrized majority rule is complementary.
@@
The conclusion of Theorem 5 is true for any gain generator ${X}$ with mean $\mu$ and any voting body formed by individualists, any number of disjoint utilitarian and majoritarian groups, and the symmetrized majority rule.
This result underlies the ‘mirrored performance’ effect discussed in Section 4.
3.5 Gains on location families of generators
For any ${X}^{0},$ which is a gain generator with mean $0,$ consider the location family of gain generators
@@ 3.5 Gains on location families of generators
For such families, the following theorem holds, which requires the symmetry of the gain distribution around its mean.
Theorem 7.
Suppose that society consists of agents with complementary strategies and employs a complementary voting rule; let the distribution of the gain generator ${X}^{\mu}$ be symmetric about $\mu.$ Then
${\mathbb{E}}\bigl{(}{\mathbf{D}}_{{\mathbf{{X}}}^{-\!\mu}}\bigr{)}={\mathbb{E}}\bigl{(}{\mathbf{D}}_{{\mathbf{{X}}}^{\mu}}\bigr{)}-\mu\bm{1},$ (8)
@@
Since the distribution of ${X}^{\mu}$ is symmetric about $\mu$, gain generators $-{X}^{\mu}$ and ${X}^{-\mu}={X}^{\mu}-2\mu$ share the same distribution. Therefore, the desired statement follows from Theorem 5. ∎
Using Lemma 2 we obtain
@@
The conclusion of Theorem 7 is true for any gain generator ${X}^{\mu}$ symmetric about $\mu$ and any voting body formed by individualists, any number of disjoint utilitarian and majoritarian groups, and the symmetrized majority rule.
4 Interpretation and discussion
@@
This rule will be referred to as the basic rule.
For any social decision rule $\psi,$ let the expected relative capital gain (ERCG) of an agent under $\psi$ be the mathematical expectation of the difference between the agent’s capital gain under $\psi$ and that under the basic rule after implementing a social decision.
Clearly, ERCG is invariant to replacing rejection with acceptance or any stochastic combination of them at $\mu=0$ in the definition of the basic rule.
For location families (7), consider the ERCG as a function of $\mu.$
@@
For any location family (7) and any agent, the expected capital gain (ECG) under the basic rule is 00 when $\mu\leq 0$ and $\mu$ when $\mu>0.$ Therefore, for any $\mu>0,$ the left-hand side ${\mathbb{E}}\bigl{(}{\mathbf{D}}_{{\mathbf{{X}}}^{-\!\mu}}\bigr{)}$ and the right-hand side ${\mathbb{E}}\bigl{(}{\mathbf{D}}_{{\mathbf{{X}}}^{\mu}}\bigr{)}-\mu\bm{1}$ of (8) are the vectors of agents’ expected relative capital gains (ERCG) under gain generators ${X}^{-\mu}$ and ${X}^{\mu}$ (which are the elements of the location family having means $-\mu$ and $\mu$), respectively. Since by Theorem 7 the left-hand side and right-hand side of (8) are equal, the ERCG of the elements of the family is an even function of $\mu.$
@@
Figure 3: Expected relative capital gain: (a) in the society of 21 individualists under majority voting (red) and voting with the optimal threshold (green); (b) in the society consisting of a group of 7 agents and 14 individualists (for comparison, the ERCG curves for homogeneous societies are added)
For the society of 21 individualists with the location family of Gaussian generators (7), the expected capital gain under simple majority rule was shown in Fig. 1. The ERCG for this family is presented in Fig. 3a. In addition to a pit of losses in unfavorable environments, this function has a mirror image of that pit in favorable environments. Decisions made by majority votes in these two areas are, on average, disadvantageous to society compared to the basic rule.
The reason for the pit of losses paradox is that in unfavorable environments, the average positive gain is smaller, while the average absolute loss is larger than in the neutral environment of the same variance (where these average values are equal). Therefore, the average difference between [the absolute values of] a gain and a loss is negative. Consequently, the proposals accepted by simple majority with a small margin of votes in favor (which is typical in an unfavorable environment) yield, on average, a total loss. Thus, in environments with $\mu<\mu_{0}<0,$ where $\mu_{0}$ is some threshold, the majority’s gain does not compensate the loss suffered by the minority. Note that this imbalance identified within the ViSE framework is relevant to practice.
The nature of the positive-area pit of losses is dual: while majorities typically approve harmful (with respect to the total capital) proposals in unfavorable environments, under favorable conditions they quite often reject proposals that provide society with positive total gains.
Thus, if for the society with parameters corresponding to Fig. 3a, environments with $\mu/\sigma<-0.254$ are considered ‘hell’ and environments with $\mu/\sigma>0.254$ are considered ‘heaven’, then simple majority rule is worse than the basic rule in both of these realms. Simple majority rule outperforms the basic rule only in a neighborhood of the neutral environment ($\mu=0$), where the ERCG peak is quite high.
@@
It is instructive to compare this with the ERCG for the societies of 21 individualists (red dotted curve) and of 21 group members (blue dotted curve). The most advantageous position is in the group of 7 agents and the worst is among the individualists complementing this group. Such a group can be considered an ‘elite’. It usually receives relatively high capital gains at the expense of individualists. It should be noted that the ERCG of the whole mixed society (gray curve) is even lower than that for the society of 21 individualists. The impact of a ‘responsible elite’ on public welfare was investigated by [43, 44]. In most cases, such an elite provides society with capital growth.
Obviously, the most beneficial strategy for the whole society is to be one cohesive group. However, in practice this is unrealistic. While a group of like-minded people is able to develop a common objective function, this is problematic for society. Therefore, non-consensus-based decisions are of primary interest. The ‘basic rule’, which outperforms the simple majority in both ‘hell’ and ‘heaven’ is extremely crude. How can taking into account voters’ opinions improve the outcome? Such a social decision rule is voting with the optimal threshold ([10, 26]). This threshold is exactly $0.5$ in the neutral environment, as well as for $\mu/\sigma\in[-0.1,0.1]$ in our example (Fig. 3a). Outside a certain interval around 00, the optimal threshold is usually (but not always) higher in unfavorable and lower in favorable environments. The ERCG curve in the case of voting with the optimal threshold clearly has no pit of losses. For the society of 21 individualists, it is shown in Fig. 3a.

### v5

@@ 3.1 Relationship between gains under opposite proposals
The following lemma establishes a connection between the implemented capital gains under proposal generators ${\mathbf{{X}}}={\mathbf{{X}}}^{\mu}$ and $\mathbf{{X}}$} (such generators will be called opposite ones) when a voting body is antisymmetric.
@@
For any gain generator X{X} with mean $\mu$ and any antisymmetric voting body, it holds that
@@
Proof.
If ${\mathbf{{x}}}\in{\mathbb{R}}^{n}$ is an arbitrary proposal generated by $\mathbf{{X}}$}, then the implemented gain vector is $\mathbf{{x}}$} with probability $I({\mathbf{{x}}},{\mathbf{{X}}})$ and 0 with probability $1-I({\mathbf{{x}}},{\mathbf{{X}}})$. Hence
${\mathbb{E}}\bigl({\mathbf{D}}_{\mathbf{{X}}}\mid{\mathbf{{X}}}={\mathbf{{x}}}\bigr)={\mathbf{{x}}}\,I({\mathbf{{x}}},{\mathbf{{X}}}).$ (2)
For the proposal generator $\mathbf{{X}}$}, we similarly have
${\mathbb{E}}\bigl({\mathbf{D}}_{-{\mathbf{{X}}}}\mid-{\mathbf{{X}}}=-{\mathbf{{x}}}\bigr)=-{\mathbf{{x}}}\,I(-{\mathbf{{x}}},-{\mathbf{{X}}}).$ (3)
Since the events $\{{\mathbf{{X}}}={\mathbf{{x}}}\}$ and $\{-{\mathbf{{X}}}=-{\mathbf{{x}}}\}$ coincide (as sets in the underlying probability space), subtracting (3) from (2) and using (1) yields
${\mathbb{E}}\bigl({\mathbf{D}}_{\mathbf{{X}}}-{\mathbf{D}}_{-{\mathbf{{X}}}}\mid{\mathbf{{X}}}={\mathbf{{x}}}\bigr)={\mathbf{{x}}}.$ (4)
Taking expectations in (4) with respect to the distribution of $\mathbf{{X}}$} gives
${\mathbb{E}}({\mathbf{D}}_{\mathbf{{X}}})-{\mathbb{E}}({\mathbf{D}}_{-{\mathbf{{X}}}})={\mathbb{E}}({\mathbf{{X}}})=\mu\bm{1},$
which proves the lemma. ∎
@@
The underlying reason for Lemma 1 is the complementarity embodied in the antisymmetry condition (1): for every fixed proposal $\mathbf{{x}}$}, the probabilities of accepting $\mathbf{{x}}$} and $\mathbf{{x}}$} by an antisymmetric voting body sum to $1$. Consequently, the conditional expectations in (2) and (3) differ precisely by $\mathbf{{x}}$}. Averaging over all proposals then yields a shift by $\mu\bm{1}$. The lemma is therefore a consequence of this reversal symmetry and relies neither on any profit-seeking properties of the voter strategies or the social decision rule nor on any distributional assumptions on X{X} other than the finiteness of its mean.
Remark 2.
An earlier version of Lemma 1 applied to societies consisting of individualists and a single group and relied on a longer proof based on the analysis of closed formulas expressing expected values. Generalizing the lemma and simplifying its proof were accomplished in two stages. In the first stage, an anonymous Reviewer suggested a simpler proof based on the observation that, under the symmetrized majority rule, the decision process with generator $\mathbf{{X}}$} is the opposite, in terms of collective decisions, of the process with generator $\mathbf{{X}}$}, viewed in reverse time. This proof was applicable to more general societies.
@@ 3.2 Complementary strategies
Consider an alternative group strategy: the group decides whether to support a proposal using the within-group majority rule. More specifically, the group members vote as individualists; let $n_{G}^{+}({\mathbf{{x}}})$ be the total number of votes they cast for proposal $\mathbf{{x}}$} in the internal group voting; finally, each group member $i$ casts $n_{i}^{+}({\mathbf{{x}}})$ votes defined by
$n_{i}^{+}({\mathbf{{x}}})=\begin{cases}1,&n_{G}^{+}({\mathbf{{x}}})>\tfrac{g}{2}\\ \tfrac{1}{2},&n_{G}^{+}({\mathbf{{x}}})=\tfrac{g}{2}\\ 0,&n_{G}^{+}({\mathbf{{x}}})<\tfrac{g}{2}\end{cases},$
where $g$ is the size of the group. The one-line form of this definition is
$\color[rgb]{0,0,0}n_{i}^{+}({\mathbf{{x}}})\;=\;\tfrac{1}{2}\big(1+\mathop{\rm sgn}\nolimits\hskip-1.1pt\big(n_{G}^{+}({\mathbf{{x}}})-\tfrac{g}{2}\big)\big).$} (5)
@@
We say that the strategy of agent $i$ is complementary if and only if
$n_{i}^{+}({\mathbf{{x}}})+n_{i}^{+}(-{\mathbf{{x}}})=1\mbox{~~for {\color[rgb]{0,0,0}all} proposals~~}{\mathbf{{x}}}\in{\mathbb{R}}^{n}.$ (6)
@@
The strategies of individualists and members of utilitarian and majoritarian groups are complementary.
Proof.
For a member of a utilitarian group, it holds that $n_{i}^{+}({\mathbf{{x}}})=\tfrac{1}{2}\hskip-1.1pt\bigl(1+\mathop{\rm sgn}\nolimits\sum_{j=1}^{g}{{x}_{j}}\bigr)$, while for a member of a majoritarian group, we have $n_{i}^{+}({\mathbf{{x}}})=\tfrac{1}{2}\hskip-1.1pt\bigl(1+\mathop{\rm sgn}\nolimits\sum_{j=1}^{g}\mathop{\rm sgn}\nolimits{{x}_{j}}\bigr).$ It is easy to verify that in both cases, Eq. (6) is valid. For an individual, this follows as a special case. ∎
3.3 Complementary social decision rules
If a stochastic voting rule that is part of a voting body uses only the number $n^{+}({\mathbf{{x}}})$ of votes cast by the society, then it can be represented by a function $V(n^{+}({\mathbf{{x}}}))$ that associates with $n^{+}({\mathbf{{x}}})$ the probability of $\mathbf{{x}}$} being accepted given that the society has cast $n^{+}({\mathbf{{x}}})$ votes for $\mathbf{{x}}$}. The value $n^{+}({\mathbf{{x}}})$, as well as its terms $n^{+}_{i}({\mathbf{{x}}})$, need not be integer.
In particular, the symmetrized majority rule has a form similar to (5):
$\color[rgb]{0,0,0}V(n^{+}({\mathbf{{x}}}))\;=\;\tfrac{1}{2}\big(1+\mathop{\rm sgn}\nolimits\hskip-1.1pt\big(n^{+}({\mathbf{{x}}})-\tfrac{n}{2}\big)\big)$}.
This rule satisfies the following condition:
$\color[rgb]{0,0,0}\mbox{For all~~}{\mathbf{{x}}}\in{\mathbb{R}}^{n},\;\>V(n^{+}({\mathbf{{x}}}))+V(n-n^{+}({\mathbf{{x}}}))=1$}. (7)
A social decision rule is also considered as part of a voting body. For such a rule that depends on $n^{+}({\mathbf{{x}}})$ and $\mathbf{{X}}$}, let $V(n^{+}({\mathbf{{x}}}),{\mathbf{{X}}})$ be the probability of accepting $\mathbf{{x}}$} generated by $\mathbf{{X}}$}. We say that this rule is complementary if and only if it satisfies the condition:
$\color[rgb]{0,0,0}\mbox{For any realization ${\mathbf{{x}}}$ of any ${\mathbf{{X}}}\in{\mathcal{X}}$,}\;\>V(n^{+}({\mathbf{{x}}}),{\mathbf{{X}}})+V(n-n^{+}({\mathbf{{x}}}),-{\mathbf{{X}}})=1$}. (8)
Complementary social decision rules act in opposite ways with respect to opposite generators.
A generalization of the symmetrized majority rule is the class of threshold social decision rules defined by
${\color[rgb]{0,0,0}V(k,{\mathbf{{X}}})\;=\;\tfrac{1}{2}(1+\mathop{\rm sgn}\nolimits(k-t({\mathbf{{X}}})))}\mbox{~~for all~~}k\in[0,n],$ (9)
where $t({\mathbf{{X}}})\in[-1,n+1]$ is the voting threshold of a rule.
Examples of such rules are the baseline rule and the optimal threshold rule discussed in Section 4.
Let us say that a threshold social decision rule is reversal-symmetric if and only if
[…]
Lemma 3.
If a threshold social decision rule is reversal-symmetric, then it is complementary.
Lemma 3 is proved by substituting (9) and (10) into the left-hand side of (8).
@@ 3.4 A class of antisymmetric voting bodies
Any voting body formed by agents with complementary strategies and a complementary social decision rule is antisymmetric.
Proof.
@@
$n^{+}({\mathbf{{x}}})+n^{+}(-{\mathbf{{x}}})=\sum_{i=1}^{n}\bigl(n_{i}^{+}({\mathbf{{x}}})+n_{i}^{+}(-{\mathbf{{x}}})\bigr)=n.$
As the social decision rule is complementary, $V(n^{+}({\mathbf{{x}}}),{\mathbf{{X}}})+V(n^{+}(-{\mathbf{{x}}}),-{\mathbf{{X}}})=1.$ Finally, the reversal symmetry, $I({\mathbf{{x}}},{\mathbf{{X}}})+I(-{\mathbf{{x}}},-{\mathbf{{X}}})=1$ for all ${\mathbf{{x}}}\in{\mathbb{R}}^{n}$, of the voting body follows from the disintegration formula (conditioning on the value of $n^{+}({\mathbf{{x}}})$), since for any $k\in{\mathbb{R}}$, $V(k,{\mathbf{{X}}})=I({\mathbf{{x}}},{\mathbf{{X}}}\mid n^{+}({\mathbf{{x}}})=k).$
@@
An example of a non-monotonic complementary voting rule is shown in Fig. 2.
By Lemma 2 and (7), Lemma 4 can be applied to societies that include several disjoint groups, along with individualists, and use the symmetrized majority rule.
@@
The sufficient conditions for antisymmetry of a voting body given in Lemma 4 are not necessary. Indeed, let the voters’ strategies support all proposals, and let the voting rule be defined by
${\color[rgb]{0,0,0}V(n^{+}({\mathbf{{x}}}))}=\begin{cases}\tfrac{1}{2},&n^{+}({\mathbf{{x}}})=n\\ 0,&n^{+}({\mathbf{{x}}})<n\end{cases}.$
Then the voting body is antisymmetric, while neither the agents’ strategies nor the voting rule are complementary. For this voting body, the conclusion of Lemma 1 takes the form ${\mathbb{E}}({\mathbf{D}}_{-{\mathbf{{X}}}})-{\mathbb{E}}({\mathbf{D}}_{\mathbf{{X}}})=-\tfrac{1}{2}\mu\bm{1}-\tfrac{1}{2}\mu\bm{1}=-\mu\bm{1}$.
This example illustrates that antisymmetry is strictly weaker than complementarity.
@@ 3.5 Complementary strategies and rules on opposite proposals
For any gain generator X{X} with mean $\mu$ and for any voting body formed by agents with complementary strategies and a complementary social decision rule,
@@
The following corollary is a consequence of Theorem 6, Lemma 2, and the fact (7) that the symmetrized majority rule is complementary.
@@
The conclusion of Theorem 6 is true for any gain generator X{X} with mean $\mu$ and any voting body formed by individualists, any number of disjoint utilitarian and majoritarian groups, and the symmetrized majority rule.
3.6 Gains on location-scale families of generators
For any ${X}^{0},$ a gain generator with mean $0,$ consider the location-scale family of gain generators
@@ 3.6 Gains on location-scale families of generators
For such families, the following corollary holds, which requires the symmetry of the gain distributions.
Corollary 8.
For any antisymmetric voting body, any location-scale family of gain generators (11) with a symmetric ${X}^{0},$ and any $\mu\in{\mathbb{R}},$
${\mathbb{E}}\bigl({\mathbf{D}}_{{\mathbf{{X}}}^{-\!\mu}}\hskip-1.1pt\bigr)={\mathbb{E}}\bigl({\mathbf{D}}_{{\mathbf{{X}}}^{\mu}}\hskip-1.1pt\bigr)-\mu\bm{1},$ (12)
@@
Since the distribution of ${X}^{\mu}$ is symmetric about $\mu$, gain generators ${X}^{\mu}$ and ${X}^{-\mu}={X}^{\mu}-2\mu$ have the same distribution. Therefore, the desired statement follows from Lemma 1. ∎
Using Corollary 8 and Lemmas 2 and 4 we obtain
@@
Under the symmetry of ${X}^{0},$ the conclusion of Corollary 8 is true, in particular, for all voting bodies formed by agents with complementary strategies and a complementary social decision rule, including voting bodies formed by individualists, any number of disjoint utilitarian and majoritarian groups, and the symmetrized majority rule.
These results underlie the “mirrored performance” effect discussed in Section 4.
4 Performance of social decision rules
@@
This rule will be referred to as the baseline rule.
For any social decision rule $\psi,$ let the expected relative capital gain (ERCG) of an agent $i$ under $\psi$ (or the performance of $\psi$ with respect to $i$) be the mathematical expectation of the difference between the $i$’s capital gain under $\psi$ and that under the baseline rule after implementing a social decision.
Clearly, ERCG is invariant to replacing rejection with acceptance or any stochastic combination of them at $\mu=0$ in the definition of the baseline rule.
For location-scale families (11), consider the ERCG as a function of $\mu.$
@@
For any family (11) and any agent, the ECG under the baseline rule is 0 when $\mu\leq 0$ and $\mu$ when $\mu>0.$ Therefore, for any $\mu>0,$ ${\mathbb{E}}\bigl({\mathbf{D}}_{{\mathbf{{X}}}^{-\!\mu}}\hskip-1.1pt\bigr)$ and ${\mathbb{E}}\bigl({\mathbf{D}}_{{\mathbf{{X}}}^{\mu}}\hskip-1.1pt\bigr)-\mu\bm{1}$ are the performance vectors of the social decision rule under quasi-opposite gain generators ${X}^{-\mu}$ and ${X}^{\mu},$ respectively. By Corollary 8, ${\mathbb{E}}\bigl({\mathbf{D}}_{{\mathbf{{X}}}^{-\!\mu}}\hskip-1.1pt\bigr)={\mathbb{E}}\bigl({\mathbf{D}}_{{\mathbf{{X}}}^{\mu}}\hskip-1.1pt\bigr)-\mu\bm{1},$ hence the performance of the social decision rule is an even function of $\mu.$
@@
Figure 3: Expected relative capital gain: (a) in the society of 21 individualists under majority voting (red) and the optimal threshold rule (green); (b) in the society consisting of a group of 7 agents and 14 individualists (for comparison, the performance curves for homogeneous societies are added)
For the society of 21 individualists with the location-scale family of Gaussian generators (11), the expected capital gain under simple majority rule was shown in Fig. 1. The performance of the simple majority rule for this family is presented in Fig. 3a. In addition to a pit of losses in hostile environments, this function has a mirror image of that pit in favorable environments. Decisions made by majority votes in these two areas are, on average, disadvantageous to society compared to the baseline rule.
The reason for the pit-of-losses paradox is that in unfavorable environments, the average positive gain is smaller, while the average absolute loss is larger than in the neutral environment of the same variance (where these average values are equal). Therefore, the average difference between the absolute values of a gain and a loss is negative. Consequently, the proposals accepted by simple majority with a small margin of votes in favor (which is typical in a hostile environment) yield, on average, a total loss. Thus, in environments with $\mu<\mu_{0}<0,$ where $\mu_{0}$ is some threshold, the majority’s gain does not compensate the loss suffered by the minority. Note that this imbalance identified within the ViSE framework is relevant to practice, and is not specific to Gaussian generators.
The nature of the positive-area pit of losses is complementary in the following sense: while majorities typically approve harmful (with respect to the total capital) proposals in hostile environments, under favorable conditions, they quite often reject proposals that provide society with positive total gains.
Thus, if for the society with parameters corresponding to Fig. 3a, environments with $\mu/\sigma<-0.254$ are considered “hell” and environments with $\mu/\sigma>0.254$ are considered “heaven”, then simple majority rule is worse than the baseline rule in both of these realms. Simple majority rule outperforms the baseline rule only in a neighborhood of the neutral environment ($\mu=0$), where the ERCG peak is quite high.
@@
It is instructive to compare this with the ERCG for the societies of 21 individualists (red dotted curve) and of 21 group members (blue dotted curve). The most advantageous position is in the group of 7 agents and the worst is among the individualists complementing this group. Such a group can be considered an “elite” in this stylized model. It usually receives relatively high capital gains at the expense of individualists. It should be noted that the ERCG of the whole mixed society (gray curve) is even lower than that for the society of 21 individualists. The impact of a “responsible elite” on public welfare was investigated by [43, 44]. In most cases, such an elite provides society with capital growth.
5 Optimal threshold rule
From a utilitarian perspective, the most beneficial strategy for the whole society would be to act as one cohesive group. However, in practice this is unrealistic. While a group of like-minded people is able to develop a common objective function, this is problematic for society. Therefore, non-consensus-based decisions are of primary interest.
The baseline rule, which outperforms the simple majority in both “hell” and “heaven” is extremely crude. How can taking into account voters’ opinions improve the outcome?
Consider the optimal threshold rule ([10, 26]). It is the threshold social decision rule whose threshold $t({\mathbf{{X}}})$ maximizes the total expected capital gain ${\mathbb{E}}({\mathbf{D}}_{\mathbf{{X}}})\bm{1}$ of the society. For simplicity, we will limit ourselves to societies of individualists.
Lemma 11.
In a society of individualists, the optimal threshold rule is reversal-symmetric.
Proof.
For any ${\mathbf{{X}}}\in{\mathcal{X}}$, $n^{+}({\mathbf{{x}}})=k+\tfrac{m}{2},$ where $k$ and $m$ are the numbers of agents with positive and zero proposed gains, respectively. Hence the condition of optimality (in the ex ante sense) of threshold $t({\mathbf{{X}}})$ (cf. [26, Theorem 2]) is $(t-\tfrac{m}{2}){\mu^{+}}+(n-t-\tfrac{m}{2})\mu^{-}=0,$ where ${\mu^{+}}={\mathbb{E}}({X}\mid{X}>0)$ and $\mu^{-}={\mathbb{E}}({X}\mid{X}<0).$ Indeed, $n^{+}({\mathbf{{x}}})>t$ (resp., $n^{+}({\mathbf{{x}}})<t$) whenever a proposal that receives $n^{+}({\mathbf{{x}}})$ votes provides a positive (resp., negative) expected gain. Therefore,
$t=\frac{n\hskip 1.1pt{|\mu^{-}|}+\tfrac{m}{2}({\mu^{+}}-{|\mu^{-}|})}{{\mu^{+}}+{|\mu^{-}|}}.$ (13)
Having in mind that $\mu^{+}_{-\hskip-0.90138pt{X}}=|\mu^{-}_{X}|$ and $|\mu^{-}_{-\hskip-0.90138pt{X}}|=\mu^{+}_{X}$, we obtain $t({\mathbf{{X}}})+t(-{\mathbf{{X}}})=n$. ∎
Now the complementarity of this rule in societies of individualists follows from Lemma 3.
Corollary 12.
For any location-scale family (11) with a symmetric ${X}^{0},$ the performance of the optimal threshold rule with respect to agents in a society of individualists is an even function of $\mu$.
Proof.
This follows from the complementarity of the optimal threshold rule in this society (Lemmas 11 and 3), the complementarity of the individualistic strategy (Lemma 2), and Corollary 10. ∎
Under the conditions of Corollary 12, the optimal threshold is $0.5$ in the neutral environment ($\mu=0$), since ${\mu^{+}}={|\mu^{-}|}$ in (13). The performance curve of the optimal threshold rule has no pit of losses. For the society of 21 individualists, it is shown in Fig. 3a. In this example, the optimal threshold is $0.5$ for $\mu/\sigma\in[-0.1,0.1]$. Outside a certain interval around 0, the optimal threshold is usually (but not always) higher in unfavorable and lower in favorable environments.